OpenStudy (anonymous):

Solve: −2(x − 3) = −6 0 −6 6 −12

OpenStudy (blurbendy):

distrubute the -2 -2x +6 = -6 -2x = -12 x = 6

OpenStudy (anonymous):

thx

OpenStudy (blurbendy):

yw